OpenStudy (anonymous):

-(3x - 2) = -4(x + 1) - 4

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm -(3x - 2) = -4(x + 1) - 4 }\) \(\large\color{midnightblue}{ \rm (-1)(3x - 2) = -4(x + 1) - 4 }\) \(\large\color{midnightblue}{ \rm -3x +2= -4x ~ -4~ - 4 }\) \(\large\color{midnightblue}{ \rm -3x +2= -4x ~ -8 }\) \(\large\color{midnightblue}{ \rm x +2= ~ -8 }\) \(\large\color{midnightblue}{ \rm x = ~ -10 }\)

OpenStudy (mathmale):

@solomonzelman: Wow! Nothing left for Emilynicoleedwards11 to do now. Is that good? The acid test is whether she can now do that by herself, thanks to your efforts.